find area of equilateral triangle of side 4a using heron’s formula. using this formula find area of an equilateral triangle whose perimeter is 540cm.
Answer by mananth(16949)

heron's formula
area of triangle = sqrt(s(s-a)(s-b)(s-c))
s= half perimeter = 4a+4a+4a /2 = 6a
Area = sqrt(6a(2a)(2a)(2a))
area = sqrt48a^4.
.
perimeter = 540
side = 180
s=270
area =sqrt(270(90)(90)(90))
=14029.26 sq cm